OpenStudy (anonymous):

Evaluate the integral of (x+3)/(x^2+9) dx

OpenStudy (anonymous):

well using fraction u get

OpenStudy (anonymous):

\[= \frac{ 1 }{2 }\ln(x^{2}+9)+\tan^{-1}(\frac{ x }{ 3 })+ C\]

OpenStudy (anonymous):

sorry i am tired i will prove that to u when i am back ok, but u can simply work it out in ur head